OpenStudy (anonymous):

Solve the system using substitution. 3x + y = 1 x - y = 3 What is the solution for x?

OpenStudy (uri):

Show your work if you tried @fukaawa

OpenStudy (immanuelv):

well first step should be to try to isolate x or y in the second equation then you can plug in the equation to the x or y in the first equation

OpenStudy (uri):

@shahd-owies Giving direct answers is against the Code of Conduct here,Please delete your answer. http://openstudy.com/code-of-conduct

OpenStudy (ikram002p):

she is new uri :P i'll kill u if u abused her again xD

OpenStudy (anonymous):

nobody ever follows the code of conduct tho lmao

OpenStudy (immanuelv):

it depends if you wanna learn or just want someone to do your homework

OpenStudy (uri):

@ikram002p I am very well aware of the fact that,that user is new...That's why I told them. Or else I could have reported :)

OpenStudy (anonymous):

3x+v=1 x-v=3 so x=3+v by substituting the value of x in the first equation 3(3+v)+v=1 9+3v+v=1 so 4v=-8 which gives that v=-2 but we knew that x=3+v which gives that x=1